While a roofer is working on a roof that slants at 40.0? above the horizontal, he accidentally nudges his 95.0N toolbox, causing it to start sliding downward, starting from rest.

If it starts 4.15m from the lower edge of the roof, how fast will the toolbox be moving just as it reaches the edge of the roof if the kinetic friction force on it is 22.0N ?
Answer #1

PE = mgh = 86 * 4.5 *sin 39 N
work done by friction
= 19 * 4.5 N
KE at the edge
= 0.5 * 86/9.81 * v^2

86 * 4.5 *sin 39 = 19 * 4.5 + 0.5 * 86/9.81 * v^2

243.55 = 85.5 + 4.38 v^2
v^2 = 36

v = 6 m/s ---answer